Grade 12Electrostatics

Two points charges placed at a distance r in air exert a force f the distance at which they exert Same force when placed in a certain medium

RANJEET Choudhary
8 Years agoGrade 12

2 Answers

Arun
8 Years ago
Dear Ranjeet
 
F (air) = (1/4 pi eo) [qQ/r^2] 
F(medium) = (1/4 pi eo*k) [qQ/R^2] 
given> F(medium) = F(air) 
(1/4 pi eo*k) [qQ/R^2] = (1/4 pi eo) [qQ/r^2] 
(1/k) [1/R^2] = [1/r^2] 
kR^2 = r^2 
R^2 = r^2/k 
R = r/ [rootk]
 
Regards
Arun (askIITians forum expert)
Rishi Sharma
6 Years ago
Dear Student,
Please find below the solution to your problem.

F (air) = (1/4 pi eo) [qQ/r^2]
F(medium) = (1/4 pi eo*k) [qQ/R^2]
given
F(medium) = F(air)
(1/4 pi eo*k) [qQ/R^2] = (1/4 pi eo) [qQ/r^2]
(1/k) [1/R^2] = [1/r^2]
kR^2 = r^2
R^2 = r^2/k
R = r/ [rootk]

Thanks and Regards